OpenStudy (anonymous):

if a clock strikes the appropriate number of time on each hour. how many times will it strikes in one day?

OpenStudy (anonymous):

i dont know?

sam (.sam.):

Let appropriate number be x, one day=24 hour, It strikes each hour, one day=24x times

OpenStudy (anonymous):

Generally the clock strikes the number of times equal to the hour hand. So if it is 1'oclock it strikes once , if it is 2'0clock it strikes twice and so on. So here in a span of 24 hours it will do a full cycle of ( 1+ 2+ 3....+12) * 2 i.e \[2 * \sum_{1}^{n} k = 2 * n* (n+1)/2\] Therefore n * (n+1) here n=12 So the clock strikes 12*13 = 156

Directrix (directrix):

From another site comes this possible solution; "Each 12 hour day it strikes 12+11+10+9+8+7+6+5+4+3+2+1 times thats 78 times but each day has 24 hours so it strikes 156 times a day. (78 X2)" http://ph.answers.yahoo.com/question/index?qid=20110304203902AAmzBTi